The positive value of a so that the co-efficient of x5 is equal to that of x15 in the expansion of is
1
2








Grade:12th Pass

1 Answers

Badiuddin askIITians.ismu Expert
13 years ago

Dear Vishwesh

general term in the series = 10Cr(x2)10-r(a/x3)r

                                                        = 10Crx20-5r(a)r

for x15   20-5r =15

              r=1

  so coefficient of this term = 10C1a

for x5    20-5r =5

             r=3

so coefficient of this term = 10C3a3

 

now equate

   10C1= 10C3a3

    a = 1/23

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin